HETAV PATEL

Grade 11,

A metal wire of linear mass density of 9.8 g/m is stretched with a tension of 10 kg-wt between two rigid supports 1 metre apart. The wire passes at its middle point between the poles of a permanent magnet, and it vibrates in resonance when carrying an alternating current of frequency n. The frequency n of the alternating source is

A metal wire of linear mass density of 9.8 g/m is stretched with a tension of 10 kg-wt between two rigid supports 1 metre apart. The wire passes at its middle point between the poles of a permanent magnet, and it vibrates in resonance when carrying an alternating current of frequency n. The frequency n of the alternating source is

Grade:12

3 Answers

anuj kumar sharma
9 years ago
the wire will vibrate with the same frequency as that of source .this can be considered as anexample of forced vibration.
T=10*9.8N=98N
M=9.8*10^-3Kg/m
Frequencyof wire,f=1/2Lroot(T/M)
                          f=1/2*1root(98/9.8*10^-3)
                           f=50Hz is the Answer.
Adeeba
5 years ago
f=1/2l√(T/m)As for1/2(1/2)/
Adeeba
5 years ago
The wire vibrate with the same frequency as that of source
f=1/2l√(T/m)
As the wire is at the middle point l=1/2
f=(2/2)/√(10*9.8)/9.8*10^-3
f=100